OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\LARGE \int ln~\sqrt[3]{x}~dx\]

OpenStudy (loser66):

take it by part, let u = ln (cuberoot x) dv =dx

OpenStudy (luigi0210):

Is my derivative right for u? \(\LARGE u=ln \sqrt[3]{x}\) and \(\LARGE dv=dx\) \(\LARGE du=\frac{1}{\sqrt[3]{x}}~dx\) and \(\LARGE v=x\) \(\LARGE (ln\sqrt[3]{x})x~-~ \int (x)(\frac{1}{\sqrt[3]{x}})~dx\) ^Evaluate that?

OpenStudy (loser66):

nope, for u du = 1/(3x) , you forgot chain rule

OpenStudy (loser66):